OpenStudy (anonymous):

28 girls went camping. There were 2 main activities: Volleyball and swimming. Fourteen girls went swimming, 5 participated in both activities, and 4 girls did neither. How many girls only played volleyball? Show work.

OpenStudy (anonymous):

28-14= 14, that is the girls that went swimming 14-5=9, that is subtracting the 5 girls that did both. 9-4=5, that is subtracting the girls that did neither. So 5 is your answer

OpenStudy (anonymous):

k so we have 2 activities: volleyball both Swimming ? 5 14 28 - 14= 14 14 - 5 = 9 - 4 = 5 < --- answer yes @s_asbury is right